The wavelength of red helium-neon laser light in air is 632.8 nm.

(a) What is its frequency?
Hz
(b) What is its wavelength in glass that has an index of refraction of 1.48?
nm
(c) What is its speed in the glass?
Mm/s
Answer #1

1)f *lamda =c (in air)

f= c/lamda = 3 *108/632.8*10-9=4.74 *1014 Hz

2)f *lamda=frequency =c/n

lamda=c/(n*f)=3 *108/(1.48*4.74 *1014) = 427.24 nm

3)frequency =c/n=3 *108 /1.48= 202.32 M m/s
Answer #2

A) f = c/lamda = 3*10^8/632.8*10^-9 = 4.74*10^14 Hz...

B) L'= L/n = 632.8/1.48 = 427.56 nm...

C) spped = L'*f = 427.56*10^-9*4.74*10^14 = 2.026*10^8 m/sec
Answer #3

a. apply V = Lf

frequency f = v/L = 3e8/632.8nm

f = 4.74*10^14 Hz

------------------------------------------------------------------

b. L = L0/n

L = 632.8nm/1.48

L =427.56 nm
-------------------------------------------

c. apply n = c/v

v = 3e8/1.48

v =2.02 *10^8 m/s

------------------------------------------------------------------------------------------------------------------------
Feel free to COMMENT on this answer if u seek furthur more CLARIFICATION. Do Rate and Encourage Me to HELP you More, HAPPY LEARNING @ CHEGG
Answer #4

a)

freq = 3x10^8/362.8x10^-9 = 4.74 x10^14 Hz

b)

from

n1/n2 = L2/L1

1/1.48 = L2/632.8

L2 = 427.56 nm

c)
speed v = c/n2 = 3x10^8 /1.48 = 2.027 x10^8 m/s
Answer #5

a. apply V = Lf

frequency f = v/L = 3e8/632.8nm

f = 4.74*10^14 Hz

------------------------------------------------------------------

b. L = L0/n

L = 632.8nm/1.48

L =427.56 nm
-------------------------------------------

c. apply n = c/v

v = 3e8/1.48

v =202.32 Mm/s

------------------------------------------------------------------------------------------------------------------------
Feel free to COMMENT on this answer if u seek furthur more CLARIFICATION. Do Rate and Encourage Me to HELP you More, HAPPY LEARNING @ CHEGG